Your starship, the Aimless Wanderer, lands on the mysterious planet Mongo. As chief scientist-engineer, you make the following measurements: a 2.50-kg stone thrown upward from the ground at 13.0 m/s returns to the ground in 5.10 s ; the circumference of Mongo at the equator is 2.80×105 km ; and there is no appreciable atmosphere on Mongo.

Part A

The starship commander, Captain Confusion, asks for the following information: what is the mass of Mongo?

Express your answer with the appropriate units.

m= ____

Part B

If the Aimless Wanderer goes into a circular orbit 40,000 km above the surface of Mongo, how many hours will it take the ship to complete one orbit?

t= ____ h
Answer #1

a = delta v / delta t

delta v = 13 m/s - (-13 m/s) = 26 m/s , delta t = 5.10 s

a = 26 / 5.10 = 5.1 m/s2

r = c / (2 pi) = ( 2.80 * 105 km) / 2 pi

r = 439823 km

a = M G / r2

M = a r2 / G M = (5.1 ) (439823 * 103m)2 / (6.67 * 10-11)

M = 1.48 * 1028 kg

---------------------------------------------------

a = g sqrt (r / (r+h))

a = (5.1) * sqrt (4.39 x105 km / (4.39 x105 km + 40000 km) )

a = 4.88 m/s2

a = v2 / r , v = (2 pi r) / P

a = ((2 pi r) / P )2 / r

P2 = (2 pi )2 r / a

P = 2 pi * sqrt (r/a)

= 2 pi * sqrt (4.39 x108 m / 4.88)

P = 16.55 hours